OpenStudy (anonymous):

what is 1.48 times 10 to the negative 5 power times 6.5 times 10 to the negative 6 power in scientific notation

OpenStudy (anonymous):

working on it

OpenStudy (anonymous):

14.8*65=962(5*6=30)932 i think

OpenStudy (anonymous):

not a 100% sure